OpenStudy (anonymous):

The sum of two numbers is 78, and their difference is 14. What are the two numbers? larger number = ____ smaller number = ____

OpenStudy (solomonzelman):

two numbers: \(\normalsize\color{royalblue}{ \rm x }\) & \(\normalsize\color{royalblue}{ \rm y }\) The sum of two numbers is 78: \(\normalsize\color{royalblue}{ \rm x+y=78 }\) their difference is 14: \(\normalsize\color{royalblue}{ \rm x-y=14 }\)

OpenStudy (solomonzelman):

So you just need to solve the system: \(\normalsize\color{royalblue}{ \rm x+y=78 }\) \(\normalsize\color{royalblue}{ \rm x-y=14 }\)

OpenStudy (solomonzelman):

when you get the numbers, tell me the answer and I will verify it for you.

OpenStudy (anonymous):

x=46 y=32

OpenStudy (solomonzelman):

yes, that is right

OpenStudy (solomonzelman):

Sorry for late reply.

OpenStudy (anonymous):

:D

OpenStudy (solomonzelman):

yw